# DS communications: Trends, Challenges, and Solutions

### v3

@@ V-B Discussions and Outlook
When it comes to DTN protocols, it has been observed from literature that some protocols in different layers may not work efficiently with one another. Hence, lack of integration in the network stack may lead to internetworking issues. In this regard, although CFDP accomplishes reliable file transmission, for example, it results in overhead when combined with BP and LTP. To address this, Kim et al. in [119] attempt to derive an optimization model to decrease the DTN network overhead for space missions caused by CFDP. However, investigating the performance of various protocols integration in the context of DS both in research and deployment is needed.
@@
The reliability of communications in DS is essential and important to ensure, but that comes at the cost of overhead. In addition to BP and its custody transfer scheme, which is the core of the DTN, a number of coding mechanisms have been adapted at the link layer. To ensure reliability even further, other coding techniques are needed even in upper layers. For example to handle the performance degradation of CFDP over long-delay and one-directional communications links, erasure coding, which helps overcoming packet loss due to DS link degradation, is suggested to be implemented [130]. Erasure coding can be used in addition to or instead of ARQ to improve data delivery as the latter has been shown to be inefficient, specially in long round-trip links [154], and is known to increase processing complexity and storage utilization unlike erasure coding which speeds up delivery process [155]. There are encouragements to conduct further research to optimize the use of erasure coding in the DTN architecture though and investigate other schemes to enhance overall DS communications performance. Even though the adoption of hop-by-hop transmission through BP is inevitably required for handling the intermittent availability of links while keeping throughput reasonable, end-to-end transmission mechanism may also be revisited along with other transport layer schemes to improve reliability avoiding permanent bundle loss in case of a node failure. One of these protocols is the deep-space transport protocol (DS-TP) which serves fast file delivery, two times faster than other conventional TPs under specific conditions [156]. However, when there is a difference in error rate estimation between a sender and a receiver, DS-TP, as discussed in [157], causes a malfunction in the re-transmission process leading to inefficient use of bandwidth. Despite the efforts made to analyze this protocol performance, much work is needed to fully understand its operation. Stream control transmission protocol (SCTP) is another TP, suggested in [158], that, in certain conditions, surpass conventional TCP and can be utilized in orbiters that operate in the same region. Some other related protocols are also proposed e.g. the seamless IP diversity based network mobility (SINEMO) suggested, in [159, 160], to be adopted by mobile entities in space while addressing the drawbacks of the network mobility (NEMO) basic support protocol (BSP), proposed to be used by NASA. These mobility protocols pose security risks, though, as in some cases the information of node locations cloud be shared in space networks. If this information is eavesdropped, connections could be exposed to major security threats. These issues need to be addressed when utilizing the mobility protocols, there are some efforts to investigate related aspects for example, authors in [161], suggested various schemes to withstand mobility protocols security threats, however, more security along with integration issues with other protocols require more investigation.
@@
CGR, due to its accuracy and efficiency, has been gaining attention recently and is shown to be the most effective routing protocol for space networks. Nonetheless, CGR faces many challenges, making it an intriguing area for researchers. Despite of the contributions and efforts done in the topic, by e.g. [153] and the references therein including [162], CGR resiliency is still a concern and subject to enhancement. In support of this claim, improvements over CGR have been shown to be possible, e.g. in [163] where earliest arrival optimal delivery ratio (EAODR) routing algorithm has been proposed. Being able to account for the earliest available and future contacts EAODR seems to offer a 12.9% delay reduction over CGR. In the same regard, addressing CGR scalability and computational complexity that may incur due to networks size increase are still research concerns. As a matter of fact, DTN scalability in general is still an open research challenge. Despite the efforts in developing DS internetworking protocols, more advanced protocols that account for the high delays, intermittent losses, network topology dynamicity, and links and entities heterogeneity while still guarantying high capacity are needed. Furthermore, DS internetworking protocol needs to account for future services such as multicast/broadcast and be efficient as satellite-to-Earth transmission to meet the demand for expanding human presence in the Solar system. Achieving redundancy by deploying more relaying systems orbiting over the different planets and moons can contribute toward achieving such a goal.
Figure 7: Next-generation IPN.
@@ VI Advances OF DSOC
Having discussed the key physical layers techniques and networking protocols along with their performance in the field of DSOC in Sections III-V, we in this section address a number of other field emerging technologies and advancements that promise to address a number of the challenges, discussed in Section II, and take DS communications to the next level. The section reviews some research and implementation efforts done in the direction of the IPN, OAM, QC, and RF/FSO systems and identifies gaps to be explored.
@@ VI-A Interplanetary Network
NASA has three operational networks that support missions: deep space network (DSN), near-Earth network (NEN), and space network (SN). DSN is the primary network that is responsible for most of the space missions communications worldwide. While DSN consists of three stations distributed around Earth separated by 120° for full coverage, NEN consists of ground stations that provide communications and tracking services from low-earth orbit to lunar orbit [164]. SN has a constellation of geosynchronous relays, which consists of 10 satellites called tracking and data relay satellite system (TDRSS) along with two ground facilities, which will undergo a complete redesigning of the architecture as arranged by the Space Network Ground Segment Sustainment (SSGS) project, to modernize the system to be capable of handling a few gigabits per second data rate capabilities [165].
@@
The goal of the IPN is to provide connectivity all around the Solar system, meeting the demand for high data volumes, including video-streaming and high quality images transmitted by the ever-increasing number of missions. The IPN will ultimately need to be accessible by the public [166]. Connecting DSN, NEN, and SN networks is the first step in deploying the IPN.
@@
Based on studies provided in [167, 168, 61], we demonstrate in Fig. 7 the vision for the IPN architecture. Currently, through RF links, Mars and the Moon elements have merely an access to the main Earth network. In the future, however, planets will have their own networks, and the IPN will become a network of networks. Each planet, as depicted in Fig. 7, will have a dedicated relay that offers a continuous connection to Earth through optical and RF-based links, referred to as a trunk links [167]. Local services between planet elements are to be made possible through other, more available, optical and RF-based, links referred to as proximity links. The orbiting satellites around Mars, Mars Atmospheric and Volatile EvolutioN (MAVEN) and Mars Odyssey, shown in the figure, are all part of the current Mars relay system. They are displayed to demonstrate Earth-Mars uplink and downlink communications. Terminal satellites may act as a link to convert optical signals to RF and vise versa; providing a significant advantage in overcoming atmospheric effects [61]. To avoid interplanetary link unavailability due to obstruction and Sun-Earth conjunction, spare interplanetary communications links through spacecrafts located at Lagrangian points are to be included [168].
@@
In addition, to implementing FSO-based relays, which are, for example, expected to offer 300 Mbps for Mars distances [167], FSO-based cubesats, which are SWaP reduced satellites and, in some cases, landers, are emerging to provide a 2 Gbps rate when thousands of them are implemented [169, 170]. The first two interplanetary cubesats launched by NASA in 2018, Mars Cube One (MarCO) probes, have successfully been deployed for navigation and monitoring [171]. The MarCO s mission, which is still operational, suggests more utilization of cubesats for future missions [172]. Lately, to improve the IPN further, it has been suggested to use asteroids as interplanetary relays [173]. It is proposed to use cubesat landers to form multiple antennas on the surface of an asteroid. Based on the numerical analysis provided in [173], an antenna of 400 m2 size provides up to 80 Mbps in Mars distances. Although cubesat is a promising technology, networking issues resulting from long communications distances can become even more severe as the number of IPN nodes becomes large. For example, when cubesats are implemented in large quantities, network protocols (NP s) like CGR, and similarly EAODR, become inefficient [174]. To mitigate that, however, rate adaptation with a bandwidth fair-sharing mechanism is suggested, for example, to enhance the process of bundle transfer and management [174].
@@ VI-B Orbital Angular Momentum
To better utilize the tremendous capacity offered by FSO, a great deal of attention is being given to exploring the potential of merging FSO with OAM, which emerges as a technology to solve the high data rate demands of future terrestrial [175, 176, 177, 178, 179, 180] and DS communications [62, 181, 182, 183]. OAM does not only have the potential to significantly enhance the spectral efficiency, but also to offer a great deal of improvement over communications energy and security.
The light that carry OAM constructs a twisted beam with a helical phase front making a singularity at the center of the beam resulting in a doughnut shape wavefront intensity profile. One of the OAM beam characteristics, commonly exploited, is the topological charge or mode number which is a positive or a negative integer that represents the number of phase twists per wavelength and the sign indicates the direction of the twist. Thanks to their orthogonality, OAM modes with infinite distinguished phase fronts provide an additional degree of freedom that results in significant capacity returns.
@@
There are various OAM generation methods that are being used in the context of DS including spiral phase plate (SPP[183], computer-generated holograms (CGH s), and spatial light modulators (SLM s). These schemes are also considered for detection. SLM, though, has shown to have a number of advantages over CGH and SPP schemes including having a higher control accuracy and resolution [182]. SPPs, in addition, has shown to have limitations when larger mode numbers are considered, accordingly, Zhang et al. in [183] suggested to use Dammann optical vortex grating (DOVG) method as it can modulate more OAM modes than the SPP and generate independent collinear OAM channel [184]. Other methods including holographic or interferometric methods are recommended to be used for obtaining high photon efficiency in DS [62, 181].
Aside from [62, 181, 182, 183], to the best of our knowledge, no other references analyzed DS OAM performance. Djordjevic et al. in [62] showed that OAM can achieve 100 Gbps in DS, however, looking at the state-of-the-art of OAM technologies, data rates can be in the range of hundreds of tera bytes per second and even higher in some cases. [177], for example, showed that multiplexing quadrature amplitude modulation (QAM) signals with OAM results in 100 Tbps for a distance of 1 m. In  [185], on the other hand, where 54.139 Gbps orthogonal frequency-division multiplexing (OFDM) 8 QAM signals over 368 WDM pol-muxed 26 OAM modes resulted in a capacity of 1.036 Pbps and spectral efficiency of 112.6 b/s/Hz. [185] has not relate such a high achievable rate to a distance though. In the context of DS, Wang et al. in [182] proposed encoding OAM states of a single photon with the deployment of relay nodes to enable a reliable and secure transmission. Analyzing the performance of the idea is missing though. In [183], weak beam intensity detection in presence of a high background noise has been realized using PC detectors and a parallel coding method evaluated with a BER as low as 1.27$\times$10-5, which is order of magnitude better than the traditional PPM.
@@
As indicated, many recent research and experimental work have, unprecedentedly, exhibited very significant results in the possibility of establishing high capacity. In addition to the general OAM implementation challenges [175], incorporating OAM in DS faces unique hurdles. We below demonstrate these hurdles and discuss attempts to address number of them.
@@
Atmospheric turbulence: OAM modes can be subject to propagation effects, which can lead to crosstalk between the modes and affect the orthogonality between OAM states. Even though AT is not of concern in space-to-space links, as stated previously, it is a significant challenge for probe-to-Earth links. When it comes to limited distance communications, mainly up to several kilometers, one can find a variety of work done for compensating the AT, see e.g [186, 187, 188, 189, 190]. In [186, 187], for example, authors have shown, through the use of deep learning techniques, 100% OAM mode recognition is possible in medium to weak turbulence levels, while between 50% [186] and 80% [188] recognition can be achieved in the case of strong turbulence. However, there are only few efforts done in the context of DS; as an example, Djordjevic et al. in [62] proposed combining the LDPC code with OAM-based modulation schemes to overcome AT effects. Considering the DS links AT mitigation methods presented in [62], which are also recommend for near Earth communications, the authors of [181] shown the possibility of the system to achieve a spectral efficiency of $N^{2}/\log_{2}(N)$ over that of LDPC-PPM modulation, where $N$ is the number of OAM modes. For the best of our knowledge, other than [181] and [62], there is no more work done addressing AT in OAM-DSOC.
@@
Lateral displacement error: Keeping an LG beam axis and its receiving entity fully aligned is difficult. For long-distance transmissions, this requirement is even more challenging. An alignment error does not only render optical power loss but also, more crucially, cause expansion of the spiral spectrum resulting in crosstalk effect and mode detection difficulty [195]. Even though it has been shown that the lateral offset can improve resiliency against eavesdropping [196], such an analysis is lacking in DS communications.
Despite the efforts done, a significant contribution yet still need to be done handling DS challenges considering the accomplishments done in limited distances moving forward to facilitating OAM implementation in DS.

### v4

@@ V-B Discussions and Outlook
When it comes to DTN protocols, it has been observed from the literature that some protocols in different layers may not work efficiently with one another. Hence, a lack of integration in the network stack may lead to internetworking issues. In this regard, although CFDP accomplishes reliable file transmission, for example, it results in overhead when combined with BP and LTP. The reliability of communications in DS is essential and important to ensure, but that comes at the cost of overhead. Kim et al. [128] attempted to derive an optimization model to decrease the DTN network overhead for space missions caused by CFDP. However, investigating the performance of various protocols integration in the context of DS both in research and deployment is needed.
@@
In addition to BP and its custody transfer scheme, which is the core of the DTN, several coding mechanisms have been adapted at the link layer. To ensure reliability even further, other coding techniques are needed even in the upper layers. For example to handle the performance degradation of CFDP over long delays and one-directional communications links, erasure coding, which helps in overcoming packet loss due to DS link degradation, is suggested to be implemented [139]. Erasure coding can be used in addition to or instead of ARQ to improve data delivery as the latter has been shown to be inefficient, especially in long round trip links [163], and is known to increase processing complexity and storage utilization unlike erasure coding which speeds up delivery [164]. There are encouragements to conduct further research to optimize the use of erasure coding in the DTN architecture and investigate other schemes to enhance overall DS communications performance. Even though the adoption of hop-by-hop transmission through BP is inevitably required to handle the intermittent availability of links while keeping throughput reasonable, end-to-end transmission mechanism may also be revisited along with other transport layer schemes to improve reliability avoiding permanent bundle loss in case of a node failure. One of these protocols is the deep space transport protocol (DS-TP) which serves fast file delivery, two times faster than other conventional TPs under specific conditions [165]. However, when there is a difference in error rate estimation between a sender and a receiver, DS-TP, as discussed in [166], causes a malfunction in the retransmission process leading to inefficient use of bandwidth. Despite the efforts made to analyze this protocol’s performance, much work is needed to fully understand its operation. Stream control transmission protocol (SCTP) is another TP, suggested in [167], that, in certain conditions, surpasses conventional TCP and can be utilized in orbiters that operate in the same region.
Some other related protocols are also proposed e.g. the seamless IP diversity based network mobility (SINEMO) suggested, in [168, 169], to be adopted by mobile entities in space while addressing the drawbacks of the network mobility (NEMO) basic support protocol (BSP), proposed to be used by NASA. These mobility protocols pose security risks, though, as in some cases the information about node locations could be shared in space networks. If this information is eavesdropped, connections could be exposed to major security threats. These issues need to be addressed when utilizing mobility protocols. There are some efforts to investigate related aspects; for example, authors in [170] suggested various schemes to withstand mobility protocols security threats. However, security and integration issues require more investigation.
@@
CGR, due to its accuracy and efficiency, has been gaining attention recently and is shown to be the most effective routing protocol for space networks. Nonetheless, CGR faces many challenges, making it an intriguing area for researchers. Despite the contributions and efforts made in the topic, by e.g. [162] and the references therein including [171], CGR resiliency is still a concern and subject to enhancement. In support of this claim, improvements over CGR have been shown to be possible, e.g. in [172] where the earliest arrival optimal delivery ratio (EAODR) routing algorithm has been proposed. Being able to account for the earliest available and future contacts, EAODR seems to offer a 12.9% delay reduction over CGR. In the same regard, addressing CGR scalability and computational complexity that may incur due to network size increase are still research concerns. As a matter of fact, DTN scalability in general is still an open research challenge. Despite the efforts made in developing DS internetworking protocols, more advanced protocols that account for high delays, intermittent losses, network topology dynamicity, and links and entities heterogeneity while still guaranteeing high capacity are needed. Furthermore, DS internetworking protocol needs to account for future services such as multicast/broadcast and ensure efficient satellite-to-Earth transmission to meet the demand for expanding human presence in the Solar system. Achieving redundancy by deploying more relaying systems orbiting over different planets and moons can contribute toward achieving such a goal.
Figure 7: Next generation IPN.
@@ VI Advances OF DSOC
Having discussed the key physical layer techniques and networking protocols along with their performance in the field of DSOC in Sections III-V, in this section we address a number of other emerging technologies and advancements that promise to address DS communications challenges and meet future field goals. The section reviews some research and implementation efforts done in the direction of the IPN, OAM, QC, and RF/FSO systems and identifies gaps to be explored.
@@ VI-A Interplanetary Network
NASA has three operational networks that support missions: deep space network (DSN), near-Earth network (NEN), and space network (SN). DSN is the primary network that is responsible for communicating with majority of space missions. While DSN consists of three stations distributed around Earth separated by 120° for full coverage, NEN consists of ground stations that provide communications and tracking services from low-Earth orbit to lunar orbit [173]. SN has a constellation of geosynchronous relays, which consists of 10 satellites called tracking and data relay satellite system (TDRSS), along with two ground facilities. SN will undergo over a redesign of its architecture, as arranged by the Space Network Ground Segment Sustainment (SSGS) project, to modernize the system to be capable of handling a few gigabits per second data rate [174].
@@
The goal of the IPN is to provide connectivity all around the Solar system meeting the demand for high-data volumes including video streaming and high-quality images transmitted by the ever-increasing number of missions. The IPN will ultimately need to be accessible by the public [175]. Connecting DSN, NEN, and SN networks is the first step in deploying the IPN.
@@
Based on [176, 177, 61], we demonstrate in Fig. 7 the vision for the IPN architecture. Currently, through RF links, Mars and Lunar entities have merely access to the main Earth network. In the future, however, planets will have their own networks, and the IPN will become a network of networks. Each planet, as depicted in Fig. 7, will have a dedicated relay that offers a continuous connection to Earth through optical and RF-based links, referred to as trunk links [176]. Local services between planet elements are to be made possible through other, more available, optical and RF-based links, referred to as proximity links. The orbiting satellites around Mars, Mars Atmospheric and Volatile EvolutioN (MAVEN) and Mars Odyssey, shown in the figure, are all part of the current Mars relay system. They are displayed to demonstrate Earth-Mars uplink and downlink communications. Terminal satellites may act as a link to convert optical signals to RF and vice versa providing a significant advantage in overcoming atmospheric effects [61]. To avoid interplanetary link unavailability due to obstruction and Sun-Earth conjunction, spare interplanetary communications links through spacecrafts located at Lagrangian points are to be included [177].
@@
In addition to implementing FSO-based relays, which are expected to offer 300 Mbps at Mars distances for example [176], Cubesats are emerging, as SWaP-reduced satellites or landers in some cases, to provide a 2 Gbps rate when thousands of them are implemented [178, 179]. The first two interplanetary Cubesats launched by NASA in 2018, Mars Cube One (MarCO) probes, have successfully been deployed for navigation and monitoring [180]. The MarCO s mission, which is still operational, suggests more utilization of Cubesats for future missions [181]. Lately, to improve the IPN further, it has been suggested to use asteroids as interplanetary relays [182]. It is proposed to use Cubesat landers to form multiple antennas on the surface of an asteroid. Based on the numerical analysis provided in [182], an antenna of 400 m2 size provides up to 80 Mbps in Mars distances. Although Cubesat is a promising technology, networking issues resulting from long communications distances can become even more severe as the number of IPN nodes becomes large. For example, when Cubesats are implemented in large quantities, network protocols like CGR, and similarly EAODR, become inefficient [183]. To mitigate that, however, rate adaptation with a bandwidth fair sharing mechanism is suggested in [183], for example, to enhance the process of bundle transfer and management.
@@ VI-B Orbital Angular Momentum
To better utilize the tremendous capacity offered by FSO, a great deal of attention is being given to exploring the potential of merging FSO with OAM, which emerges as a technology to solve the high data rate demands of future terrestrial [184, 185, 186, 187, 188, 189] and DS communications [62, 190, 191, 192, 193]. OAM does not only have the potential to significantly enhance spectral efficiency, but also to offer a great deal of improvement over communications energy and security. The light that carries OAM constructs a twisted beam with a helical phase front making a singularity at the center of the beam resulting in a doughnut-like shape wavefront intensity profile. One of the OAM beam characteristics commonly exploited is the topological charge or mode number, which is a positive or negative integer that represents the number of phase twists per wavelength and the sign indicates the direction of the twist. Thanks to their orthogonality, OAM modes with infinite distinguished phase fronts provide an additional degree of freedom that results in significant capacity returns. There are various OAM generation methods that are being used in the context of DS including spatial light modulators (SLM s), spiral phase plate (SPP s) [192], computer-generated holograms (CGH s). These schemes are also used for detection. SLM, which could be followed by a charge-coupled device (CCD) camera when used as a detector [191], has shown to have several advantages over CGH and SPP schemes including having a higher-control accuracy and resolution [191]. SPPs, in addition, have been shown to have limitations when larger-mode numbers are considered. Accordingly, Zhang et al. [192] suggested using the Dammann optical vortex grating (DOVG) method as it can modulate more OAM modes than the SPP and generate independent collinear OAM channels [194]. Other methods including holographic or interferometric methods are recommended to be used for obtaining high-photon efficiency in DS [62, 190].
There is a handful of work done exploring the performance of implementing OAM in DS. From the capacity aspect, Djordjevic et al. [62] showed that DS OAM link can achieve a rate of 100 Gbps, based on LDPC-coded multidimensional OAM modulation. By employing 10 Gbps of data on 10 OAM dimensions, that could be provided by using PPM for instance [190], the aggregated rate of 100 Gbps has shown to be achievable. Higher rates are accomplished by increasing the number of states used. However, scaling to higher states comes at the cost of increasing the crosstalk between the modes, thereby lowering the performance of BER. Looking at the OAM state-of-the-art technologies, data rates can be in the range of hundreds of terabytes per second and even higher in some cases. Huang et al. [186], for example, showed that multiplexing quadrature amplitude modulation (QAM) signals with OAM results in 100 Tbps for a distance of 1 m. Wang et al. [195], on the other hand, demonstrated that 54.139 Gbps orthogonal frequency division multiplexing (OFDM) 8 QAM signals over 368 WDM pol-muxed 26 OAM modes resulted in a capacity of 1.036 Pbps and spectral efficiency of 112.6 b/s/Hz. The authors in [195] have not related such a high-achievable rate to a distance though.
Wang et al. [191] proposed encoding OAM states of a single photon for DS communications. They also discussed that the number of states does not affect channel efficiency and hence, crosstalk is not of concern when all communications entities are beyond the Earth’s atmosphere. Higher-order states have been avoided in the analysis provided in [191] though, as they are more susceptible to diffraction. Furthermore, Wang et al. [191] suggested that the deployment of relay nodes within OAM DS links to facilitate a reliable and secure transmission. In [192], on the other hand, a weak beam intensity detection in the presence of high-background noise, measured in counts per second (cps), has been realized using a single photon PC detector and parallel coding. The proposed technique, which is referred to as OAM-pair PPM coding, is suggested as an efficient method suitable for long distances including DS. Utilizing this scheme, each OAM-pair achieves $(\log_{2}(M)+1)$ bits per pulse resulting in ${n}/{2}$ multiple of that total pulse density, when $n$ even number of OAM modes are transmitted over PPM signal with order $M$. According to their analysis, the authors demonstrated that 12×12$\times$12 × increment in the capacity of PPM can be achieved. Regarding the BER performance, OAM-pair PPM achieves orders of magnitude improvement over the traditional PPM, as Fig. 8 demonstrates.
Figure 8: Parallel coded OAM performance [192, Fig. 3].
@@
Many recent research and experimental work have, unprecedentedly, exhibited very significant results in the possibility of establishing high capacity. Nonetheless, aside from [62, 190, 191, 192], to the best of our knowledge, no more investigations have been dedicated in the context of DS. In addition to the general OAM implementation challenges [184], incorporating OAM in DS faces unique hurdles. We below demonstrate these hurdles and discuss the attempts to address a number of them.
@@
Atmospheric turbulence: OAM modes can be subject to propagation effects, which can lead to crosstalk between modes and affect the orthogonality between OAM states. Even though AT is not of concern in space-to-space links, as stated previously, it is a significant challenge for probe-to-Earth links. When it comes to limited-distance communications, mainly up to several kilometers, one can find a variety of work done for compensating AT, see e.g. [196, 197, 198, 199, 200]. In [196, 197], for example, authors have shown, through the use of deep learning techniques, that 100% OAM mode recognition is possible in medium to weak turbulence levels, while between 50% [196] and 80% [198] recognition can be achieved in the case of strong turbulence. Zhao et al. [196] used an image classification based diffractive neural network to detect 10 OAM modes under AT. They indicated that at a significant level of turbulence, detection accuracy drops very low due to the turbulence-induced spiral spreading that causes crosstalk between the OAM modes posing detection limitations. Furthermore, Hao et al. [198], used a conventional neural network to detect 16 OAM modes, under various turbulence levels based on image processing while taking into account the complexity and computation time to ensure practicality. It is relevant to consider AT for DS links due to the use of ground entities in many scenarios. However, the success of deep learning and image classification in addressing AT in the context of DS has never been explored, to the best of our knowledge. Other methods like correction codes could also be implemented to mitigate AT. In the context of DS, Djordjevic et al. [62] proposed combining the LDPC code with OAM-based modulation schemes to overcome AT effects. Considering the DS links AT mitigation methods presented in [62], which are also recommended for near-Earth communications, the authors of [190] shown the possibility of the system to withstand strong turbulence and achieve a spectral efficiency of $n^{2}/\log_{2}(n)$ times that of LDPC-PPM modulation for $n$ number of modes. Noting that [62, 190] are the only efforts attempted in regard to taking AT into consideration for DS.
@@
The good news, though, that could make reception possibly reasonable at ultra-long distances, despite the huge spatial spread, is the fact that OAM modes have been experimentally proven to be detected even if they are partially received [202, 203, 204, 205]. Modes detection through receiving a very small portion of the whole circumference of an electromagnetic wave carrying OAM has been reported by [206]. Surprisingly, partially receiving a beam can lead to a reduction in crosstalk effects, surpassing the performance of full area receivers in some cases. Furthermore, according to [205], a limited size full area receiver that does not capture the maximum of the OAM mode intensity profile is outperformed by a partial receiver positioned at the maximum intensity of the beam. As indicated in [203], for two overlapping beams, two detectors receiving three quarters of the sectional area at an SNR $30$ dB achieve $12$ b/s/Hz higher-spectral efficiency than a complete area receiver. Deng et al. [204], in addition, proposed an approach that resulted in a 46.2% reduction in crosstalk intensity, achieving as low as 7.71% with four non-adjacent OAM states and no more than 22% for two adjacent modes, indicating that closer mode numbers induce more crosstalk. Despite the OAM partial reception promises shown in limited communications distances, however, as beams are largely spread in DS, the minimum portion of a beam that is needed for a mode to be recognized could be relatively large. To the best of our knowledge, the partial signal detection of OAM beams at DS distances, has never yet been investigated. Analyzing its feasibility using conventional methods as well as artificial intelligence based schemes would be of interest.
Lateral displacement error: Keeping an LG beam axis and its receiving entity fully aligned is difficult. For long-distance transmissions, this requirement is even more challenging. An alignment error does not only render optical power loss, but also, more crucially, cause expansion of the spiral spectrum resulting in a crosstalk effect and mode detection difficulty [207]. Even though it has been shown that the lateral offset can improve resiliency against eavesdropping [208], and can be mitigated with the use of lateral correction methods [209], such an analysis is lacking in DS communications.
Despite the efforts made, a significant contribution still needed to handle DS challenges considering the accomplishments made in limited distances moving forward to facilitating OAM implementation in DS.